### v2

# S2-Diffusion: Generalizing from Instance-level to Category-level Skills in Robot Manipulation

Abstract
Recent advances in skill learning has propelled robot manipulation to new heights by enabling it to learn complex manipulation tasks from a practical number of demonstrations. However, these skills are often limited to the particular action, object, and environment instances that are shown in the training data, and have trouble transferring to other instances of the same category. In this work we present an open-vocabulary Spatial-Semantic Diffusion policy (S2-Diffusion) which enables generalization from instance-level training data to category-level, enabling skills to be transferable between instances of the same category. We show that functional aspects of skills can be captured via a promptable semantic module combined with a spatial representation. We further propose leveraging depth estimation networks to allow the use of only a single RGB camera. Our approach is evaluated and compared on a diverse number of robot manipulation tasks, both in simulation and in the real world. Our results show that S2-Diffusion is invariant to changes in category-irrelevant factors as well as enables satisfying performance on other instances within the same category, even if it was not trained on that specific instance. Full videos of all real-world experiments are available in the supplementary material.
I INTRODUCTION
Imitation learning (IL) [1, 2] has shown potential in enabling robotic manipulation in challenging real-world scenarios by learning complex skills from human demonstrations. Still, existing IL methods often struggle to generalize beyond the specific training environments from which the demonstrations are derived. This is an important obstacle as each new environment requires labor-intensive data collection, model fine-tuning, and retraining to adapt the learned policies.
For humans, transferring knowledge between tasks and skills, such as transferring the scooping skill from rice to cereals, is rather straightforward. Scooping rice or cereals may be considered as different instances of the same task for current IL methods. The ability to generalize over such instances is still a challenge and requires rather advanced spatial-semantic understanding [3]. The ability to transfer and generalize over instances removes the necessity for extensive training and also allows for assessing what type of instances one can transfer over - for example, scooping ice cream may be very different from scooping granular material such as cereals or rice. Thus, granular materials may be seen as the same category as rice and cereals, while ice cream is an instance of another category which requires a very different policy when executing the scooping task.
Figure 1: Our open-vocabulary Spatial-Semantic Diffusion policy (S2-Diffusion) not only efficiently completes the task at hand but also enables the generalization of the same skill across diverse contexts and task variations. For instance, a rice-scooping policy can be directly transferred to scoop cereal without any fine-tuning.
Large pre-trained Visual-Language-Action (VLA) models [4, 5, 6] generalize simple skills such as pick-and-place over a wide range of environments and objects. However, more complex non-prehensile manipulation such as scooping, mug-flipping, cooking shrimp or opening a bottle with a bottle opener stay elusive for such large general models. Recent imitation learning approaches address learning of instances of challenging manipulation tasks  [1, 2, 7, 8, 9, 10, 11] but the integration with semantic knowledge in highly-complex manipulation tasks remains a challenge [9]. The aforementioned methods often rely on raw perceptual features and environmental conditions, limiting their applicability to the instances observed during the training. Training a skill via imitation learning, such as a diffusion policy [9], relies on expert demonstrations that often do not cover several instances of the same task. We show that when trained only on a particular instance of a task - such as wiping red scribbles from a whiteboard - the skill fails to transfer already when the scribbles are now green, even if the required action and environment for wiping is exactly the same. This is because the policy did not learn a whiteboard-wiping category skill but a single instance of this category namely red-whiteboard-wiping.
Motivated by the above challenges, we present a novel approach that integrates spatial prediction with semantic segmentation features from large pre-trained models [12, 13] to generalize from expert demonstrations on a single instance of a task to its category-level skill – such as wiping different color scribbles from the whiteboard or scooping different granular material from bowl to bowl. Our method uses a single RGB camera view combined with the proprioceptive information of the robot. As shown in Fig. 1, we extract semantic information using the prompted foundation model [14] and combine it with a depth estimation foundation model [15] to obtain Spatial-Semantic observations for the visuomotor diffusion policy learning framework. This allows for invariance to task-irrelevant factors such as background and object textures as well as the capability to generalize from instance-level to category-level skills. Our key contributions are threefold:
  1. 1.
    We propose to endow visuomotor diffusion policies with spatial-semantic understanding to enable generalization from instance-level to category-level skills in robot manipulation tasks.
  2. 2.
    We introduce an efficient representation of the spatial-semantic information via a combination of vision foundation models. The overall framework is real-time viable and requires only a single RGB camera and the robot’s proprioceptive observations.
  3. 3.
    Our extensive experiments evaluate the method on a set of robotics manipulation tasks in simulation and the real world, demonstrating the ability to learn generalizable and effective robotic manipulation policies.
All real-world experimental videos ($174$) are available in the supplementary material.
II RELATED WORK
II-A Visual Feature Based Imitation Learning
In recent studies, visual imitation learning methods [16, 17, 18, 19] have shown great potential in diverse robot manipulation tasks. The recent work of Zhu et al. [3] demonstrates the potential of object-based representations for manipulation by using a conventional closed-set object detector. In [20], the authors validate that large-scale video generative pre-training is able to effectively benefit visual robot manipulation learning. However, in their works, the finetuned policy is not demonstrated to generalize to novel instances of the task. 3D vision based imitation learning [21, 22, 23, 24] has shown to improve performance but typically requires multiple posed views with RGB-D sensors to reconstruct the scene. Another work [25] also ulitizes pretrained VLMs to provide scene representation for robot tasks, but their approach needs to construct a 3D neural feature field by scanning part of the scene. In our paper, we propose a visuomotor imitation learning method that takes advantage of object-aware semantic representation to focus on task-relevant objects for training a visuomotor policy.
Recently, Wang et al. [26] introduce a novel framework that enhances generalization capabilities by incorporating explicit spatial and semantic information into policy learning. This method uses multi-view RGB-D observations to generate 3D descriptor fields and then constructs semantic fields that highlight task-relevant parts of objects (e.g., a knife handle or soda can tab). By combining 3D geometric and semantic information, this approach achieves significant improvements in generalization at the cost of requiring multiple RGB-D cameras.
II-B Vision Language Models in Robotics
Vision-Language Models (VLMs), trained on vast amounts of internet data, have found widespread use in robotic applications [27, 28, 29]. Some works [30, 31, 32, 26] have explored leveraging 2D foundation vision models to construct open-vocabulary 3D representations. These approaches involve utilizing visual features from pretrained foundation models like CLIP [27] or SAM [33] to scene representations that benefit learning downstream robot policy. The most recent works [12, 34, 35] demonstrate superior open-vocabulary classification ability which succeed in generalizing to unseen classes. Pretrained foundation models are also utilized by either being prompted or fine-tuned to generate high-level task plans [36, 31] and to shape reward functions [37, 38, 39] for reinforcement learning. In [40], the authors introduce a multimodal diffusion transformer policy for solving long-horizon tasks. In contrast, our approach concentrates on improving generalization ability for open-vocabulary manipulation policy. MOKA [41] addresses robotic manipulation by using VLMs to mark open-world keypoint affordances. The core idea behind both MOKA and our method is similar, that is guiding the policy to focus on task-relevant objects. However, our method utilizes a more generalizable semantic representation and also depth information for the entire scene.
II-C Vision Language Actions in Robotics
Large policies or Vision-Language Actions (VLAs), pretrained on a combination of Internet-scale data and diverse robot demonstrations, is another promising research direction for generalizable policy learning in robotics. Recent works [42, 5, 43, 44] propose to co-finetune robot trajectory data alongside vision-language data, demonstrating impressive generalization capabilities with novel objects and instructions. However, these are complex and computationally costly attempts at general manipulation that while providing encouraging results at task reasoning have so far not been able to compete with imitation-based skill learning on non-prehensile and contact-rich manipulation tasks. To enhance generalization ability, [45] further proposes a cross-embodied robot policy trained on the diverse robot dataset to control robots with varying observation and action types. For comparison, our method focuses on policy generalization across different tasks by concentrating solely on the segmentation masks of target objects.
III Towards Generalizable Robotic Skills
We start with a question: "What does it mean for a skill to be generalizable?". Commonly, "mug flipping" trained on a specific mug with distinct appearance and geometry will not generalize to flipping other mugs. Instead, the learned skill remains instance-specific, as neural network-based methods optimize the training loss without explicit interpretability of the learned features. Consequently, if these (self-selected) features are absent in another mug instance, the skill will fail to perform effectively.
One approach to mitigate this limitation is to explicitly learn features that are shared across all instances within a category. For example, a "mug" could be characterized by shared geometric features, such as a cylindrical body and a handle. Works such as ShapeNet [46] provide large-scale repositories of 3D shapes categorized into object classes, enabling methods that are able to learn category-level features that can be generalized over individual instances [47]. These approaches enable generalization across objects within the same class while preserving distinctions between categories.
This object-centric approach has proven effective for tasks involving well-categorized objects like mugs, shoes, pens, or soda cans. However, tasks like wiping or scooping present unique challenges that cannot be addressed solely by these methods. In this work, we follow a more action-centric approach, that considers the semantic functionality of the skill rather than specific object categories.
In this context, the concept of affordances [48] provides a useful lens to interpret our action-centric approach. Affordances, which describe the potential actions that an object or environment can support, emphasize functionality over rigid categorization. While our work focuses on learning lower-level skills and contact-rich tasks, which are difficult to reduce to affordances, it shares a similar philosophy by focusing on the semantic functionality of skills rather than specific object categories.
III-A Functional and Semantic Abstractions
When considering robotic skills, it is essential to move beyond instance-specific implementations and emphasize a leveled abstraction of skills. This facilitates generalization across instances and categories, aligning with the notion that skills can share functional similarities despite differing in object appearances, materials, or environmental contexts. In Fig. 2 we show an example of such a skill levels - depicting the functional goals on the top, able to be achieved with different category-level skills and performed on instances of that category.
Figure 2: Skill level example for flipping and scooping. The functional goal is stated as an abstract action on the top row, second row shows examples of category-level skills that fall under this functional similarity. The bottom row shows Instances of the respective categories.
As the example in Fig. 2 shows, the skill flipping represents a generic functional task: "the act of inverting an object". Within this abstraction, mug flipping is a category-level skill that applies to any object classified as a "mug". However, when trained specifically on a single mug (e.g., "a blue cylindrical mug with a narrow handle"), the skill becomes an instance-specific skill, which may not generalize to mugs with varying shapes, textures, or weights. Similarly, the skill scooping provides another illustrative example. At its core, scooping refers to the functional action of transferring a granular or semi-solid material from one location to another. Category-level skill: Bowl-to-bowl scooping captures the general action of transferring materials between two containers, encompassing a wide range of granular and semi-solid materials. Instance-level skills: Tasks such as rice scooping or cereal scooping are more specific, as they have different visual appearances as well as grain size and density.
In this work, we define an instance-level skill as a skill learned from specific demonstrations that are tightly coupled to a particular object, material, or environmental setup. For example, a rice-bowl-to-bowl scooping policy trained exclusively on a dataset containing only rice would constitute an instance-level skill, as the robot may fail to generalize when attempting to scoop a different granular material, such as cereal or sand. On the other hand, a category-level skill is a more abstract representation of a skill that remains valid across multiple instances within a category. For example, bowl-to-bowl scooping encompasses a variety of granular materials, such as rice, cereal, or lentils, enabling the robot to execute the task reliably across these different instances. Category-level skills incorporate a more generalized functional understanding, allowing the robot to adapt to new but related scenarios without explicit retraining.
In practice, all training data - in the real world - will come from instance-specific setups for a category-level skill. One way to achieve generalization is simple to present the imitation learning method with a wide range of instances in order to cover the whole category and achieve generalization. This approach, however, has the drawback of being very labor-intensive and therefore costly when done in the real world. We address this shortcoming by focusing on Spatial-Semantic features that can be extracted from a single instance but are still informative enough so that generalization can happen. In other words when we learn a visuomotor diffuson policy with Spatial-Semantic observations - instead of RGB - we can generalize from instance-level to its category-level skill.
Figure 3: S2-Diffusion Architecture. The architecture is composed of three components: a pretrained semantic segmentation model Grounded-SAM2 [14], a pretrained depth prediction model DepthAnythingV2 [13] and a U-Net denoising diffusion policy [9]. We design an object-aware spatial-semantic representation that is leveraged for denoising probabilistic model.
IV PROBLEM FORMULATION
Let $\Psi$ be the space of RGB images. Consider a manipulation task given the language instruction of the target object $\mathcal{L}$ (e.g., "coffee mug"). Our goal is to learn a generalizable imitation learning policy by leveraging spatial-semantic representation from pretrained Vision-Language models.
We consider a robotic task as a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho,\gamma)$ of states, actions, transition probability, reward, initial state distribution, and discount factor. We assume access to a dataset $D=\{\tau_{i}\}_{i}^{N}$ of $N$ demonstrated expert trajectories $\tau_{i}=\{(s_{0},a_{0}),...,(s_{T_{i}},a_{T_{i}})\}$ for the task. Our method uses a denoising diffusion process [9] to learn imitation policy $\pi$ from $D$. To generalize from individual instances to unseen instances from the same category, we propose to build a spatial-semantic representation. In this paper, $\mathcal{S}$ is the state space composed of visual spatial-semantic representation $f_{v}$ and robot proprioception states $q$. $\mathcal{A}$ is the action space of robot end-effector commands. We aim to learn a policy $\pi_{\theta}(a|s):\mathcal{S}\rightarrow\mathcal{A}$ with parameters $\theta$ that predicts action $a$ according to current state $s$ for an MDP by leveraging the prior experience contained in the dataset $D$.
V METHOD
Our objective is to develop an open-vocabulary spatial-semantic visuomotor policy that can generalize from an individual instance to other unseen instances resulting in a category-level skill. We propose open-vocabulary Spatial-Semantic Diffusion policy (S2-Diffusion), an approach that leverages three main components in policy learning: a semantic segmentation model, a depth prediction model and a diffusion policy shown in Fig. 3. The policy is trained with demonstrations from expert teleoperation, using only RGB images and robot proprioception as the state, and end-effetor velocities as the commanded actions respectively.
V-A Diffusion for Robot Skill Learning
Following previous works [9, 49], we formulate the visuomotor policy as a conditional Denoising Diffusion Probabilistic Model (DDPM) [50, 51]. Starting from random action $a^{K}$ sampled from Gaussian noise, the diffusion probabilistic model $\epsilon_{\theta}$ performs $K$ iterations of denoising. This process gradually produces a sequence of actions with decreasing noise levels, $a^{K},a^{k-1},...,a^{0}$, until the noise-free action $a^{0}$. Each action denoising iteration is described as:
$a^{k-1}=\alpha_{k}(a_{k}-\gamma_{k}\epsilon_{\theta}(a^{k},o,k))+\sigma_{k}\mathcal{N}(0,I),\$ (1)
where $o$ is the observation for the policy. $\alpha_{k}$, $\gamma_{k}$ and $\sigma_{k}$ are referred as noise schedule for each iteration $k$, and $\mathcal{N}(0,I)$ is the Gaussian noise added to the action.
To learn the action predicting model $\pi_{\theta}$, we randomly sample the robot action $a^{0}$ from the demonstration dataset $D$ and add the noise $\epsilon^{k}$ for a random iteration $k$. The training loss for the diffusion model is formulated as:
$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{0}+\epsilon^{k},o,k)),\$ (2)
where we use an action sampling approach rather than a noise prediction model to enhance the generation of high-dimensional actions.
V-B Open-Vocabulary Spatial-Semantic Representation
We aim to design spatial-semantic-aware representation that is leveraged for the observation of the above denoising probabilistic model. We utilize two pretrained VLMs, Grounded-SAM2 [14] and DepthAnythingV2 [13], for open-vocabulary semantic segmentation and depth map estimation respectively. We use Grounded-SAM2 model to perform zero-shot semantic segmentation, leveraging CLIP-based [27] mask classification to segment unseen classes. We combine the extracted features to construct a spatial-semantic representation that is leveraged as the input for the visuomotor diffusion policy.
We utilize Grounded-SAM2 to segment an image into a set of semantic masks $(z_{1},z_{2},...,z_{n})$ from visual observations based on text descriptions. We apply pixel-wise maximum pooling for each pixel location $(i,j)$ across all segmentation masks:
$z_{f}(i,j)=\max_{s=1}^{n}z_{s}(i,j),$ (3)
where $z_{s}(i,j)$ represents the pixel value at position $(i,j)$ in the $s$-th mask. In this way, we combine multiple segmentation masks into a single mask $z_{f}$ where each pixel represents the most confident prediction from the set of masks.
Input : Semantic query $l$, image observation $o_{t}$, robot state $q$, Grounded-SAM2 model $C_{1}$, DepthAnythingV2 model $C_{2}$
1 Collect demonstrated trajectories $D=\{\tau_{i}\}_{i}^{N}$
2 for epoch n=1,N do
3       Sample raw image $o_{t}$ and robot action $a^{0}$
4       Obtain spatial-semantic representation
5                 $z_{s}=C_{1}(o_{t},l)$, $z_{d}=C_{2}(o_{t})$
6                 $z_{f}(i,j)=\max_{s=1}^{n}z_{s}(i,j)$
7                 $f_{v}=\text{ResNet}(z_{f}\|z_{d})$
8
9      Add Gaussian noise $\epsilon^{k}\sim\mathcal{N}(0,I)$ for step $k$
10                 $a^{k}=a^{0}+\epsilon^{k}$
11
12      Train the policy
13                 $\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{k},k,f_{v},q))$
14 end for
15
return the trained policy $\pi_{\theta}(a_{t}|o_{t},q,k)$
Algorithm 1 Learning Open-Vocabulary Spatial-Semantic Diffusion Policy
To improve the spatial reasoning of the imitation policy, we propose to incorporate depth map of the task space into its semantic representation. Specifically, the input image is also processed separately with a pretrained 335M-parameter DepthAnythingV2 model [13] that shows promising performance in fine-grained details. The Depth-Anything model predicts depth estimates relative to the input RGB observations rather than absolute values. This can result in inconsistencies and inaccuracies over extended tasks or manipulation horizons. To address this, we normalize the depth maps during both training and evaluation, ensuring consistency across diverse scenes and improving the model’s robustness and reliability.
The resulting semantic and spatial feature vectors are concatenated along the channel dimension to form the spatial-semantic representation. Upon generating the spatial-semantic representation, we leverage it as input for a visuomotor diffusion policy [9].
Figure 4: Simulated Tasks. We perform evaluations on five single-stage tasks from a large-scale simulation framework RoboCasa [52]: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda. RoboCasa provides 50 expert demonstrations for each of the manipulation tasks. In the demonstration datasets for each task, we assume variations in the color and shape of the target object, as well as differences in the background environment.
TABLE I: Success Rate of Simulation Experiments
Method Tasks
ServeMug CloseDoor TurnOnMicrowave TurnOffFaucet MoveSoda
BC-RNN 0.00 0.02 0.00 0.26 0.02
BC-Transformer 0.14 0.58 0.78 0.26 0.08
Diffusion Policy 0.12 0.42 0.34 0.32 0.16
S2-Diffusion (Ours) 0.72 0.78 0.84 0.70 0.82
V-C Learning Semantic Diffusion Policy
To effectively utilize our open-vocabulary spatial-semantic representations, we adopt the CNN-based diffusion policy architecture [9] as our decision-making backbone. We use Denoising Diffusion Implicit Models (DDIM) [53] as the noise scheduler. In our paper, we denote the observation with $s_{t}$ that is composed of spatial-semantic feature $f_{v}$ and robot proprioceptive state $q$. The training loss in Equation 2 is defined as:
$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(\alpha_{k}a^{0}+\beta_{k}\epsilon^{k},k,f_{v},q)),\$ (4)
where $\alpha_{k}$ and $\beta_{k}$ are used for noise schedule of each step. The learning process for our S2-Diffusion method is described in Algorithm 1. We combine features inferred by VLMs Grounded-SAM2 and DepthAnythingV2 to construct the spatial-semantic representation for the action denoising model. By conditioning on the tuple $(a^{k},k,f_{v},q)$, the denoising model learns to predict the clean action by using mean square error loss for action supervision.
VI EVALUATION
The goals of all our experiments are three-fold:
  1. 1.
    to evaluate and compare the performance of our method on challenging robotic manipulation tasks;
  2. 2.
    to validate that generalization from instance-level to category-level skill is achieved in a real-world setting;
  3. 3.
    to perform ablations in order to investigate the role of the semantic and spatial components.
We first describe the experiments performed in the simulation, followed by the experiments on physical hardware.
VI-A Simulation Experiments
Experiment Setup. For our simulation experiments, we take advantage of a recent open-sourced large-scale simulation environment, RoboCasa [52], which provides expert demonstrations for diverse everyday tasks. We evaluate our method and the baselines in a set of single-stage tasks as shown in Fig. 4: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda using the provided $50$ expert demonstrations from RoboCasa.
Baselines. We compare our method with three baseline methods in simulation: 1) BC-RNN: a behavior cloning method with recurrent neural network implementation; 2) BC-Transformer: a behavior cloning imitation learning method with transformer architecture as the policy implementation [54]; 3) Diffusion Policy: the image-based diffusion policy of [9]. For real-world experiments, we compare our method against the diffusion policy. In all our experiments, we train for $500$ epochs on an NVIDIA RTX 4090 GPU. We use a batch size of $64$ and a learning rate of $1\text{e}{-4}$.
Simulation Results. We take advantage of the open-sourced implementation for the baseline methods from RoboMimic [54]. In the demonstration datasets for each task, we assume variations set by RoboCasa in the color and shape of the target object, as well as differences in the background environment. To test the performance of our method and the baselines, we evaluate $50$ trials for each task and the corresponding success rates are listed in Table I, demonstrating the superior performance of the proposed S2-Diffusion approach.
Due to the challenging variations in target object and background seen within the RoboCasa dataset, classical imitation learning policies struggle to solve the specific single-stage manipulation task. While the baseline methods—BC-RNN, BC-Transformer, and Diffusion Policy—show relatively poor performance, particularly on tasks like ServeMug, CloseDoor, and MoveSoda, S2-Diffusion consistently achieves the highest success rates. For instance, S2-Diffusion outperforms the other methods with a success rate of 0.72 on ServeMug and 0.78 on CloseDoor, while the best-performing baseline, BC-Transformer, achieves only 0.14 and 0.58, respectively. On other tasks like TurnOnMicrowave and MoveSoda, S2-Diffusion maintains a strong performance with success rates of 0.84 and 0.82. These results highlight the effectiveness of S2-Diffusion in improving task success across a diverse range of manipulation tasks. Similar to findings in RoboCasa [52], the image-based Diffusion Policy is sensitive to in object color and background. In contrast, by leveraging spatial-semantic features our method shows better robustness and generality compared to state-of-the-art alternatives. We attribute this to the rich visual representation of combining semantic information and depth estimation for the workspace scene.
VI-B Real-World Experiments
Experiment setup. We evaluate our S2-Diffusion method on two real-world category tasks: whiteboard-wiping and bowl-to-bowl scooping. We collected $40$ and $60$ expert demonstrations for the red-whiteboard-wiping and rice-bowl-to-bowl scooping tasks instances respectively. The demonstrations were obtained by teleoperating a $7$-DOF Franka Panda manipulator using the Quest2ROS [55] Oculus app. A single camera providing RGB observation was mounted on the end-effector as well as task-specific tools - such as a sponge for the whiteboard wiping and a spoon for the scooping task. The teleoperation setup is shown in Fig. 5. The language prompts for obtaining the semantic observations were "handwriting. sponge." for the whiteboard wiping task and "rice. bowl." for the bowl-to-bowl scooping task.
Figure 5: Real-world setup for the wiping task showing the teleoperation controller held by the operator on the right.
Baselines and Ablations. We compare our S2-Diffusion method with the visuomotor diffusion method from [9] which uses RGB and proprioceptive information as a baseline, as well as perform a number of ablations on the scooping task - namely a version of our method that only has access to semantic observations (Semantic-Diffusion) and one that only observes the spatial observations (Spatial-Diffusion).
Figure 6: Comparison of our S2-Diffusion and the baseline on two real-world environments: whiteboard wiping and bowl-to-bowl scooping. S2-Diffusion and the baseline are trained on red-whiteboard-wiping dataset and rice-bowl-to-bowl-scooping dataset respectively, then evaluated on the known instances and transferred to unseen instances of the two tasks. Note that for choco-cereal-btb-scooping, hearts-cereal-btb-scooping, mixed-cereal-btb-scooping, and green-whiteboard-wiping the baseline diffusion policy shows $0\%$ success rate.
Real-World Results. The real-world experimental results are reported in Fig. 6. We evaluate our method and the baseline first on the seen environment (red-whiteboard-wiping and rice-bowl-to-bowl-scooping) and evaluate the transferability on other instances of the same category - namely black-whiteboard- and green-whiteboard-wiping and choco-cereal-, hearts-cereal-, mixed-cereal-bowl-to-bowl-scooping.
Whiteboard-wiping: The task involves wiping scribbles of different colors (red, black, green) from a whiteboard using the end-effector-mounted wet sponge. The results are shown in Fig. 6. We define success as the robot completely removing the targeted handwriting using a sponge within $15$ seconds. For the task instances that are covered by the training data (red-whiteboard-wiping), the baseline is able to succeed in all nine trials, leading to a $100\%$ success rate, however, once the policy is deployed on different instances of the task such as black-whiteboard-wiping the performance deteriorates to $4/9$ ($44\%$) and for the green-whiteboard-wiping task even to $0\%$.
This clearly shows how even small changes in the RGB observations such as changing the color of the scribbles can lead to great deterioration and even complete failure of the skill on this instance - underscoring that the skill learning was indeed on the instance-level and not category-level. On the other hand, our S2-Diffusion method has consistent performances across all tasks. These results highlight the ability of our S2-Diffusion to learn category-level skills for handling novel tasks without requiring additional training or fine-tuning, a capability that is essential for real-world robotic applications.
Bowl-to-bowl-scooping: The scooping task is to scoop granular materials from one bowl to another. The amount of material successfully scooped into the target bowl is measured in grams. Task success is defined as the policy’s ability to scoop at least $3$ grams of material into the target bowl in under $30$ seconds. In Fig. 6 the success rates are reported over $10$ trials. All real-world experimental videos are available on the project website.
For the in-distribution rice-bowl-to-bowl-scooping task, our S2-Diffusion and baseline achieve success rates of $1.0$ and $0.8$, respectively. However, when transferring the policy to three unseen tasks—–choco-cereal-btb-scooping, hearts-cereal-btb-scooping, and mixed-cereal-btb-scooping—–the baseline diffusion policy fails entirely, with $0.0$ success rate, while our S2-Diffusion policy maintains a high success rate of approximately $0.8$ by changing the semantic prompt to "cereal. bowl." for all three instances. This demonstrates that our method effectively generalizes from individual instances to unseen other instances of the same category.
We can see frames of the baseline diffusion policy and our method in Fig. 7, where the baseline can only succeed on the instance-level task it was trained on and fails to scoop out cereal as the visual observations are too different. Our method on the other hand extracts the semantic mask from the prompt as well as the spatial information via the synthetic depth observation. This leads to the successful execution of the scooping task and to the learning of a successful category-level skill trained on individual instances only using a single RGB observation (same as the baseline) as the original input.
Figure 7: Frames of the baseline diffusion policy and our S2-Diffusion method for real-world scooping tasks. We show the corresponding semantic mask and synthetic depth for each frame. The baseline can only succeed on the instance-level task it was trained on. Our method succeeds in learning a category-level skill trained on individual instances only using a single RGB observation.
VI-C Ablation Evaluation
To investigate the significance of integrating spatial-semantic representation, we conduct additional experiments by training the policy using only the semantic representation (Semantic-Diffusion) and only the depth map estimation (Spatial-Diffusion) on the rice and mixed-cereal-bowl-to-bowl-scooping tasks. The results of these experiments are presented in Fig. 8. The full model S2-Diffusion, which integrates both modalities, achieves the highest success rates, significantly outperforming the ablated versions both on the seen task and also the novel task.
When the depth information is excluded, the performance of the policy drops drastically, with success rates falling to $0.4$ and $0.5$ for the seen and novel tasks, respectively. This significant decline underscores the critical role that spatial information plays in successfully executing 3D tasks. In contrast, when the mask (semantic representation) is removed, the performance also declines but to a lesser extent compared to the removal of depth information. While the policy’s ability to generalize across tasks is still reasonable in this case, it clearly benefits from the combined use of both modalities.
From these results, we conclude that both semantic representation and depth estimation are indispensable for learning a generalizable skill trained on individual instances. The integration of spatial-semantic features enhances the robustness of the policy, enabling better performance across both familiar and novel environments, and ultimately leading to more effective task execution in 3D spaces.
Figure 8: Ablation study of Semantic and Spatial-diffusion only on a seen task instance (left) and an unseen task instance (right). We show how the semantic component and spatial component contribute to the performance of the policy. S2-Diffusion outperforms models trained with only semantic or spatial observations.
VII Limitations
As Fig. 2 shows, the functional goal of flipping or scooping is discretized in different category-level skills such as bowl-to-bowl and pile-to-container scooping. While our method does generalize from instance-level tasks to category-level tasks, in its current form it is not able to generalize well from one instance-level skill to an instance of a different category task. i.e. when only training on rice-bowl-to-bowl the performance will degrade on a sand-pile-to-container task.
While Fig. 2 makes the ordering of skills into instance and category levels look straightforward, there are some corner cases that would need to be addressed if this view was expanded into a more complex ontology. For one, the functional goal can depend on more context than what is shown in our example. For instance, flipping is presented in our example as a way to turn over an object, but it could also be interpreted as an action the robot should perform with its own body. Furthermore, given how current learning-from-demonstration techniques are implemented, the category bowl-to-bowl scooping might require additional qualifiers to accurately encompass the correct instances. For example, if we were to generalize this skill to ice cream bowl-to-bowl scooping, the policy would fail, as the forces required are completely different from those observed in the training instances. Ice cream, especially when frozen, requires significantly more force to be scooped in the first place. However, as the ice cream begins to melt, and the required forces become more similar to those seen during training, there is a point at which our category skill bowl-to-bowl scooping would likely work. We leave exploring such nuances to future work.
Another limitation of our work is its dependency on the performance of pre-trained vision foundation models. Naturally if the semantic or spatial estimate is poor, the resulting category-level skill may generalize poorly. To address this, we plan to investigate fine-tuning VLMs in conjunction with the downstream robot policy network in future research.
VIII CONCLUSION
In this work, we have introduced S2-Diffusion, an open-vocabulary spatial-semantic diffusion policy that enables generalization from instance-level training data to category-level skills in robotic manipulation. Our method integrates semantic understanding with spatial representations, leveraging vision foundation models to learn a policy that is invariant to task-irrelevant visual changes and generalizes across different instances of the same category. This allows robots to transfer learned skills beyond their training data without requiring additional fine-tuning or retraining.
Through extensive simulations and real-world evaluations, we demonstrated that S2-Diffusion outperforms the baselines. In particular we showed that:
  • Spatial-Semantic representations enhance generalization – By combining spatial information with semantic segmentation, our method enables robots to focus on task-relevant features, improving instance-to-category transfer.
  • Efficient real-time execution using only a single RGB camera – Unlike approaches that require multi-view setups or additional depth sensors, S2-Diffusion extracts meaningful scene representations from a single RGB image, making it practical for real-world deployment.
  • Category-level generalization – Our evaluations on a number of robotic manipulation tasks, including real-world wiping and scooping, demonstrate that the method successfully generalizes across unseen instances within the same category, achieving high performance where baseline policies fail.
In summary, S2-Diffusion policies are less constrained by their training distribution and constitute a step toward the ability of humans to generalize skills across variations in objects, materials, and environments instances.

### v3

@@
 Our results show that S2-Diffusion is invariant to changes in category-irrelevant factors as well as enables satisfying performance on other instances within the same category, even if it was not trained on that specific instance.
-Full videos of all real-world experiments are available in the supplementary material.
+Project website: https://s2-diffusion.github.io.
+Index Terms:
+Imitation Learning, Learning from Demonstration, Deep Learning in Grasping and Manipulation.
 I Introduction
@@
 Thus, granular materials may be seen as the same category as rice and cereals, while ice cream is an instance of another category which requires a very different policy when executing the scooping task.
-Figure 1: Our open-vocabulary Spatial-Semantic Diffusion policy (S2-Diffusion) not only efficiently completes the task at hand but also enables the generalization of the same skill across diverse contexts and task variations.
-For instance, a rice-scooping policy can be directly transferred to scoop cereal without any fine-tuning.
+Figure 1: Our Spatial-Semantic Diffusion policy (S2-Diffusion) not only efficiently completes the task at hand but also enables the generalization of the same skill across diverse contexts and task variations.
 Large pre-trained Visual-Language-Action (VLA) models [4, 5, 6] generalize simple skills such as pick-and-place over a wide range of environments and objects.
 However, more complex non-prehensile manipulation such as scooping, mug-flipping, cooking shrimp or opening a bottle with a bottle opener stay elusive for such large general models.
-Recent imitation learning approaches address learning of instances of challenging manipulation tasks  [1, 2, 7, 8, 9, 10, 11] but the integration with semantic knowledge in highly-complex manipulation tasks remains a challenge [9].
+Recent imitation learning approaches address learning of instances of challenging manipulation tasks  [2, 7, 8, 9, 10] but the integration with semantic knowledge in highly-complex manipulation tasks remains a challenge.
 The aforementioned methods often rely on raw perceptual features and environmental conditions, limiting their applicability to the instances observed during the training.
-Training a skill via imitation learning, such as a diffusion policy [9], relies on expert demonstrations that often do not cover several instances of the same task.
+Training a skill via imitation learning, such as a diffusion policy [2], depends on expert demonstrations that often do not cover several instances of the same task.
 We show that when trained only on a particular instance of a task - such as wiping red scribbles from a whiteboard - the skill fails to transfer already when the scribbles are now green, even if the required action and environment for wiping is exactly the same.
@@
 This allows for invariance to task-irrelevant factors such as background and object textures as well as the capability to generalize from instance-level to category-level skills.
-Our key contributions are threefold:
-  1. 1.
-    We propose to endow visuomotor diffusion policies with spatial-semantic understanding to enable generalization from instance-level to category-level skills in robot manipulation tasks.
-  2. 2.
-    We introduce an efficient representation of the spatial-semantic information via a combination of vision foundation models.
+Our key contributions are threefold: i) We propose to endow visuomotor diffusion policies with spatial-semantic understanding to enable generalization from instance-level to category-level skills in robot manipulation tasks.
+ii) We introduce an efficient representation of the spatial-semantic information via a combination of vision foundation models.
 The overall framework is real-time viable and requires only a single RGB camera and the robot’s proprioceptive observations.
-  3. 3.
-    Our extensive experiments evaluate the method on a set of robotics manipulation tasks in simulation and the real world, demonstrating the ability to learn generalizable and effective robotic manipulation policies.
-All real-world experimental videos ($174$) are available in the supplementary material.
+iii) Our extensive experiments evaluate the method on a set of robotics manipulation tasks in simulation and the real world, demonstrating the ability to learn generalizable and effective robotic manipulation policies.
+All real-world experimental videos ($246$) are on the project website.
 II RELATED WORK
-II-A Visual Feature Based Imitation Learning
-In recent studies, visual imitation learning methods [16, 17, 18, 19] have shown great potential in diverse robot manipulation tasks.
-The recent work of Zhu et al.
-[3] demonstrates the potential of object-based representations for manipulation by using a conventional closed-set object detector.
-In [20], the authors validate that large-scale video generative pre-training is able to effectively benefit visual robot manipulation learning.
-However, in their works, the finetuned policy is not demonstrated to generalize to novel instances of the task.
-3D vision based imitation learning [21, 22, 23, 24] has shown to improve performance but typically requires multiple posed views with RGB-D sensors to reconstruct the scene.
-Another work [25] also ulitizes pretrained VLMs to provide scene representation for robot tasks, but their approach needs to construct a 3D neural feature field by scanning part of the scene.
-In our paper, we propose a visuomotor imitation learning method that takes advantage of object-aware semantic representation to focus on task-relevant objects for training a visuomotor policy.
-Recently, Wang et al.
-[26] introduce a novel framework that enhances generalization capabilities by incorporating explicit spatial and semantic information into policy learning.
-This method uses multi-view RGB-D observations to generate 3D descriptor fields and then constructs semantic fields that highlight task-relevant parts of objects (e.g., a knife handle or soda can tab).
-By combining 3D geometric and semantic information, this approach achieves significant improvements in generalization at the cost of requiring multiple RGB-D cameras.
-II-B Vision Language Models in Robotics
-Vision-Language Models (VLMs), trained on vast amounts of internet data, have found widespread use in robotic applications [27, 28, 29].
-Some works [30, 31, 32, 26] have explored leveraging 2D foundation vision models to construct open-vocabulary 3D representations.
-These approaches involve utilizing visual features from pretrained foundation models like CLIP [27] or SAM [33] to scene representations that benefit learning downstream robot policy.
-The most recent works [12, 34, 35] demonstrate superior open-vocabulary classification ability which succeed in generalizing to unseen classes.
-Pretrained foundation models are also utilized by either being prompted or fine-tuned to generate high-level task plans [36, 31] and to shape reward functions [37, 38, 39] for reinforcement learning.
-In [40], the authors introduce a multimodal diffusion transformer policy for solving long-horizon tasks.
-In contrast, our approach concentrates on improving generalization ability for open-vocabulary manipulation policy.
-MOKA [41] addresses robotic manipulation by using VLMs to mark open-world keypoint affordances.
-The core idea behind both MOKA and our method is similar, that is guiding the policy to focus on task-relevant objects.
-However, our method utilizes a more generalizable semantic representation and also depth information for the entire scene.
-II-C Vision Language Actions in Robotics
-Large policies or Vision-Language Actions (VLAs), pretrained on a combination of Internet-scale data and diverse robot demonstrations, is another promising research direction for generalizable policy learning in robotics.
-Recent works [42, 5, 43, 44] propose to co-finetune robot trajectory data alongside vision-language data, demonstrating impressive generalization capabilities with novel objects and instructions.
-However, these are complex and computationally costly attempts at general manipulation that while providing encouraging results at task reasoning have so far not been able to compete with imitation-based skill learning on non-prehensile and contact-rich manipulation tasks.
-To enhance generalization ability, [45] further proposes a cross-embodied robot policy trained on the diverse robot dataset to control robots with varying observation and action types.
-For comparison, our method focuses on policy generalization across different tasks by concentrating solely on the segmentation masks of target objects.
+Visual Feature Based Imitation Learning.
+Visual imitation learning methods [15, 16] have shown strong potential in diverse robot manipulation tasks.
+While Zhu et al.
+[3] and Wu et al.
+[17] highlight the benefits of object-centric and large-scale video pretraining, they do not demonstrate generalization to novel task instances.
+3D vision-based methods [18, 19, 20] improve generalization but require costly multi-view RGB-D setups or scene scanning.
+In contrast, our method uses a semantic-aware representation from a single RGB image.
+A concurrent work, SAM2Act [21], introduces a multi-view transformer-based robotics policy for improved visual feature representation and multitask generalization, with SAM2Act+ further incorporating a memory-based architecture for episodic recall in spatial memory-dependent manipulation tasks.
+Closest to our work, Wang et al.
+[22] integrate explicit spatial-semantic information into policy learning, but rely on multi-view 3D descriptor fields, whereas we achieve similar benefits without extra hardware.
+Vision-Language Models in Robotics.
+Vision-Language Models (VLMs) trained on large internet corpora [23, 24] have become prominent in robotics.
+Recent works [25, 22] leverage 2D foundation models like CLIP [23] and SAM [26] to construct open-vocabulary 3D representations.
+Newer models [11, 27] improve open-world generalization.
+VLMs are also employed to generate high-level task plans [28, 29] or reward functions [30, 31].
+Reuss et al.
+[32] introduce multimodal diffusion policies for long-horizon tasks.
+MOKA [33] uses keypoint affordances for manipulation, but we enhance generalization further by incorporating both semantic and depth information.
+Vision-Language Actions in Robotics.
+Vision-Language Actions (VLAs) pretrained on large-scale data [34, 5, 35] show strong task reasoning but remain computationally intensive and struggle with fine-grained manipulation.
+Cross-embodied policies [36] attempt to scale generalization across robots.
+Instead, our method focuses on compact, efficient policies by leveraging open-vocabulary segmentation masks for policy generalization across manipulation tasks.
+Figure 2: Skill abstraction hierarchy for flipping and scooping tasks.
 III Towards Generalizable Robotic Skills
-We start with a question: "What does it mean for a skill to be generalizable?".
-Commonly, "mug flipping" trained on a specific mug with distinct appearance and geometry will not generalize to flipping other mugs.
-Instead, the learned skill remains instance-specific, as neural network-based methods optimize the training loss without explicit interpretability of the learned features.
-Consequently, if these (self-selected) features are absent in another mug instance, the skill will fail to perform effectively.
-One approach to mitigate this limitation is to explicitly learn features that are shared across all instances within a category.
-For example, a "mug" could be characterized by shared geometric features, such as a cylindrical body and a handle.
-Works such as ShapeNet [46] provide large-scale repositories of 3D shapes categorized into object classes, enabling methods that are able to learn category-level features that can be generalized over individual instances [47].
-These approaches enable generalization across objects within the same class while preserving distinctions between categories.
-This object-centric approach has proven effective for tasks involving well-categorized objects like mugs, shoes, pens, or soda cans.
-However, tasks like wiping or scooping present unique challenges that cannot be addressed solely by these methods.
-In this work, we follow a more action-centric approach, that considers the semantic functionality of the skill rather than specific object categories.
-In this context, the concept of affordances [48] provides a useful lens to interpret our action-centric approach.
-Affordances, which describe the potential actions that an object or environment can support, emphasize functionality over rigid categorization.
-While our work focuses on learning lower-level skills and contact-rich tasks, which are difficult to reduce to affordances, it shares a similar philosophy by focusing on the semantic functionality of skills rather than specific object categories.
-III-A Functional and Semantic Abstractions
-When considering robotic skills, it is essential to move beyond instance-specific implementations and emphasize a leveled abstraction of skills.
-This facilitates generalization across instances and categories, aligning with the notion that skills can share functional similarities despite differing in object appearances, materials, or environmental contexts.
-In Fig.
-2 we show an example of such a skill levels - depicting the functional goals on the top, able to be achieved with different category-level skills and performed on instances of that category.
-Figure 2: Skill level example for flipping and scooping.
-The functional goal is stated as an abstract action on the top row, second row shows examples of category-level skills that fall under this functional similarity.
-The bottom row shows Instances of the respective categories.
-As the example in Fig.
-2 shows, the skill flipping represents a generic functional task: "the act of inverting an object".
-Within this abstraction, mug flipping is a category-level skill that applies to any object classified as a "mug".
-However, when trained specifically on a single mug (e.g., "a blue cylindrical mug with a narrow handle"), the skill becomes an instance-specific skill, which may not generalize to mugs with varying shapes, textures, or weights.
-Similarly, the skill scooping provides another illustrative example.
-At its core, scooping refers to the functional action of transferring a granular or semi-solid material from one location to another.
-Category-level skill: Bowl-to-bowl scooping captures the general action of transferring materials between two containers, encompassing a wide range of granular and semi-solid materials.
-Instance-level skills: Tasks such as rice scooping or cereal scooping are more specific, as they have different visual appearances as well as grain size and density.
-In this work, we define an instance-level skill as a skill learned from specific demonstrations that are tightly coupled to a particular object, material, or environmental setup.
-For example, a rice-bowl-to-bowl scooping policy trained exclusively on a dataset containing only rice would constitute an instance-level skill, as the robot may fail to generalize when attempting to scoop a different granular material, such as cereal or sand.
-On the other hand, a category-level skill is a more abstract representation of a skill that remains valid across multiple instances within a category.
-For example, bowl-to-bowl scooping encompasses a variety of granular materials, such as rice, cereal, or lentils, enabling the robot to execute the task reliably across these different instances.
-Category-level skills incorporate a more generalized functional understanding, allowing the robot to adapt to new but related scenarios without explicit retraining.
-In practice, all training data - in the real world - will come from instance-specific setups for a category-level skill.
-One way to achieve generalization is simple to present the imitation learning method with a wide range of instances in order to cover the whole category and achieve generalization.
-This approach, however, has the drawback of being very labor-intensive and therefore costly when done in the real world.
-We address this shortcoming by focusing on Spatial-Semantic features that can be extracted from a single instance but are still informative enough so that generalization can happen.
-In other words when we learn a visuomotor diffuson policy with Spatial-Semantic observations - instead of RGB - we can generalize from instance-level to its category-level skill.
+What does it mean for a skill to be generalizable?
+Commonly, skills like "mug flipping" learned on a specific instance do not transfer to new instances with different appearance or geometry.
+Neural policies often overfit to instance-specific features, limiting their applicability.
+One solution is to explicitly learn features shared across a category.
+For instance, "mugs" can be characterized by a cylindrical body and handle.
+Datasets like ShapeNet [37] facilitate learning such category-level features.Object-centric representations work well for defined object classes (e.g., mugs, shoes), however, they struggle with tasks involving materials or actions, such as wiping or scooping.
+Instead, we advocate an action-centric perspective, focusing on semantic functionality over object identity.
 Figure 3: S2-Diffusion Architecture.
@@
 We design an object-aware spatial-semantic representation that is leveraged for denoising probabilistic model.
+Skills should be abstracted beyond specific instances.
+As illustrated in Fig.
+2, a functional goal like scooping encompasses various category-level skills and their specific instances.
+For example, scooping represents transferring granular or semi-solid materials.
+A category-level skill like bowl-to-bowl scooping covers materials such as rice, cereal, or lentils, whereas instance-level policies (e.g., rice scooping) often fail to generalize.
+In real-world robotics, demonstrations naturally reflect instance-level setups.
+Generalization could be attempted by collecting demonstrations across many instances, but this is costly and labor-intensive.
+Instead, we propose leveraging Spatial-Semantic features extracted from a single instance.
+By replacing raw RGB inputs with spatial-semantic observations, we enable visuomotor diffusion policies to generalize from instance-level training to category-level execution.
 IV PROBLEM FORMULATION
-Let $\Psi$ be the space of RGB images.
-Consider a manipulation task given the language instruction of the target object $\mathcal{L}$ (e.g., "coffee mug").
 Our goal is to learn a generalizable imitation learning policy by leveraging spatial-semantic representation from pretrained Vision-Language models.
-We consider a robotic task as a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho,\gamma)$ of states, actions, transition probability, reward, initial state distribution, and discount factor.
-We assume access to a dataset $D=\{\tau_{i}\}_{i}^{N}$ of $N$ demonstrated expert trajectories $\tau_{i}=\{(s_{0},a_{0}),...,(s_{T_{i}},a_{T_{i}})\}$ for the task.
+We assume access to a dataset $D=\{\tau_{i}\}_{i}^{N}$ of $N$ demonstrated expert trajectories $\tau_{i}=\{(o_{0},a_{0}),...,(o_{T_{i}},a_{T_{i}})\}$ for the task.
 Our method uses a denoising diffusion process [2] to learn imitation policy $\pi$ from $D$.
 To generalize from individual instances to unseen instances from the same category, we propose to build a spatial-semantic representation.
-In this paper, $\mathcal{S}$ is the state space composed of visual spatial-semantic representation $f_{v}$ and robot proprioception states $q$.
+$\mathcal{O}$ is the observation space composed of visual spatial-semantic representation $f_{v}$ and robot proprioception states $q$.
 $\mathcal{A}$ is the action space of robot end-effector commands.
-We aim to learn a policy $\pi_{\theta}(a|s):\mathcal{S}\rightarrow\mathcal{A}$ with parameters $\theta$ that predicts action $a$ according to current state $s$ for an MDP by leveraging the prior experience contained in the dataset $D$.
+We aim to learn a policy $\pi_{\theta}(a|o):\mathcal{O}\rightarrow\mathcal{A}$ with parameters $\theta$ that predicts action $a$ according to current observation $o$ by leveraging the prior experience contained in the dataset $D$.
 V METHOD
@@
 Each action denoising iteration is described as:
-$a^{k-1}=\alpha_{k}(a_{k}-\gamma_{k}\epsilon_{\theta}(a^{k},o,k))+\sigma_{k}\mathcal{N}(0,I),\$ (1)
+$a^{k-1}=\alpha_{k}(a_{k}-\gamma_{k}\epsilon_{\theta}(a^{k},o,k))+\sigma_{k}\mathcal{N}(0,I),\\$ (1)
 where $o$ is the observation for the policy.
@@
 The training loss for the diffusion model is formulated as:
-$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{0}+\epsilon^{k},o,k)),\$ (2)
+$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{0}+\epsilon^{k},o,k)),\\$ (2)
 where we use an action sampling approach rather than a noise prediction model to enhance the generation of high-dimensional actions.
+Input : Semantic query $l$, image observation $o_{t}$, robot state $q$, Grounded-SAM2 model $C_{1}$, DepthAnythingV2 model $C_{2}$
+1 Collect demonstrated trajectories $D=\{\tau_{i}\}_{i}^{N}$
+2 for epoch n=1,N do
+3   Sample raw image $o_{t}$ and robot action $a^{0}$
+4   Obtain spatial-semantic representation
+5        $z_{s}=C_{1}(o_{t},l)$, $z_{d}=C_{2}(o_{t})$
+6        $z_{f}(i,j)=\max_{s=1}^{n}z_{s}(i,j)$
+7        $z=z_{f}\oplus z_{d}$
+8        $f_{v}=\text{ResNet}(z)$
+9
+10  Add Gaussian noise $\epsilon^{k}\sim\mathcal{N}(0,I)$ for step $k$
+11        $a^{k}=a^{0}+\epsilon^{k}$
+12
+13  Train the policy
+14        $\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{k},k,f_{v},q))$
+15 end for
+return the trained policy $\pi_{\theta}(a_{t}|o_{t},q,k)$
+Algorithm 1 Learning Open-Vocabulary Spatial-Semantic Diffusion Policy
 V-B Open-Vocabulary Spatial-Semantic Representation
@@
 In this way, we combine multiple segmentation masks into a single mask $z_{f}$ where each pixel represents the most confident prediction from the set of masks.
-Input : Semantic query $l$, image observation $o_{t}$, robot state $q$, Grounded-SAM2 model $C_{1}$, DepthAnythingV2 model $C_{2}$
-1 Collect demonstrated trajectories $D=\{\tau_{i}\}_{i}^{N}$
-2 for epoch n=1,N do
-3       Sample raw image $o_{t}$ and robot action $a^{0}$
-4       Obtain spatial-semantic representation
-5                 $z_{s}=C_{1}(o_{t},l)$, $z_{d}=C_{2}(o_{t})$
-6                 $z_{f}(i,j)=\max_{s=1}^{n}z_{s}(i,j)$
-7                 $f_{v}=\text{ResNet}(z_{f}\|z_{d})$
-8
-9      Add Gaussian noise $\epsilon^{k}\sim\mathcal{N}(0,I)$ for step $k$
-10                 $a^{k}=a^{0}+\epsilon^{k}$
-11
-12      Train the policy
-13                 $\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{k},k,f_{v},q))$
-14 end for
-15
-return the trained policy $\pi_{\theta}(a_{t}|o_{t},q,k)$
-Algorithm 1 Learning Open-Vocabulary Spatial-Semantic Diffusion Policy
 To improve the spatial reasoning of the imitation policy, we propose to incorporate depth map of the task space into its semantic representation.
 Specifically, the input image is also processed separately with a pretrained 335M-parameter DepthAnythingV2 model [12] that shows promising performance in fine-grained details.
-The Depth-Anything model predicts depth estimates relative to the input RGB observations rather than absolute values.
+The model predicts depth estimates $z_{d}$ relative to the input RGB observations rather than absolute values.
 This can result in inconsistencies and inaccuracies over extended tasks or manipulation horizons.
 To address this, we normalize the depth maps during both training and evaluation, ensuring consistency across diverse scenes and improving the model’s robustness and reliability.
-The resulting semantic and spatial feature vectors are concatenated along the channel dimension to form the spatial-semantic representation.
+The resulting semantic and spatial feature vectors are concatenated along the channel dimension to form the spatial-semantic representation $z=z_{f}\oplus z_{d}$, where $z\in\mathbb{R}^{(C_{s}+C_{d})\times H\times W}$ with $C_{s},C_{d}$ representing the channel dimensions of the semantic and depth features.
 Upon generating the spatial-semantic representation, we leverage it as input for a visuomotor diffusion policy [2].
 Figure 4: Simulated Tasks.
-We perform evaluations on five single-stage tasks from a large-scale simulation framework RoboCasa [52]: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda.
-RoboCasa provides 50 expert demonstrations for each of the manipulation tasks.
-In the demonstration datasets for each task, we assume variations in the color and shape of the target object, as well as differences in the background environment.
-TABLE I: Success Rate of Simulation Experiments
+We perform evaluations on six single-stage tasks from a large-scale simulation framework RoboCasa [41]: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda, TurnOnStove, and two tasks in SAPIEN simulator: HangMug, InsertPencil.
+TABLE I: Success Rate of Simulation Experiments in RoboCasa
 Method Tasks
-ServeMug CloseDoor TurnOnMicrowave TurnOffFaucet MoveSoda
-BC-RNN 0.00 0.02 0.00 0.26 0.02
-BC-Transformer 0.14 0.58 0.78 0.26 0.08
-Diffusion Policy 0.12 0.42 0.34 0.32 0.16
-S2-Diffusion (Ours) 0.72 0.78 0.84 0.70 0.82
+ServeMug CloseDoor TurnOnMicrowave TurnOffFaucet MoveSoda TurnOnStove
+BC-RNN 0.02 [0.01, 0.07] 0.03 [0.01, 0.08] 0.05 [0.02, 0.11] 0.27 [0.19, 0.36] 0.06 [0.03, 0.12] 0.26 [0.18, 0.35]
+BC-Transformer 0.16 [0.10, 0.24] 0.55 [0.45, 0.64] 0.70 [0.60, 0.78] 0.31 [0.23, 0.41] 0.10 [0.06, 0.17] 0.34 [0.25, 0.44]
+Diffusion Policy 0.15 [0.09, 0.23] 0.52 [0.42, 0.62] 0.41 [0.32, 0.51] 0.38 [0.29, 0.48] 0.22 [0.15, 0.31] 0.50 [0.40, 0.60]
+S2-BC-Transformer 0.54 [0.44, 0.63] 0.77 [0.68, 0.84] 0.72 [0.63, 0.80] 0.55 [0.45, 0.64] 0.44 [0.35, 0.54] 0.60 [0.50, 0.69]
+S2-Diffusion (Ours) 0.75 [0.66, 0.82] 0.80 [0.71, 0.87] 0.80 [0.71, 0.87] 0.77 [0.68, 0.84] 0.83 [0.74, 0.89] 0.88 [0.80, 0.93]
+TABLE II: Simulation Experiments in SAPIEN
+Method InsertPencil HangMug
+Seen Unseen Seen Unseen
+GenDP 0.95 [0.76, 0.99] 0.85 [0.64, 0.95] 0.90 [0.70, 0.97] 0.80 [0.58, 0.92]
+S2-Diffusion (Ours) 0.90 [0.70, 0.97] 0.90 [0.70, 0.97] 0.85 [0.64, 0.95] 0.90 [0.70, 0.97]
 V-C Learning Semantic Diffusion Policy
@@
 We use Denoising Diffusion Implicit Models (DDIM) [42] as the noise scheduler.
-In our paper, we denote the observation with $s_{t}$ that is composed of spatial-semantic feature $f_{v}$ and robot proprioceptive state $q$.
+In our paper, the observation is composed of spatial-semantic feature $f_{v}$ and robot proprioceptive state $q$.
 The training loss in Equation 2 is defined as:
-$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(\alpha_{k}a^{0}+\beta_{k}\epsilon^{k},k,f_{v},q)),\$ (4)
+$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(\alpha_{k}a^{0}+\beta_{k}\epsilon^{k},k,f_{v},q)),\\$ (4)
 where $\alpha_{k}$ and $\beta_{k}$ are used for noise schedule of each step.
@@
 For our simulation experiments, we take advantage of a recent open-sourced large-scale simulation environment, RoboCasa [41], which provides expert demonstrations for diverse everyday tasks.
-We evaluate our method and the baselines in a set of single-stage tasks as shown in Fig.
-4: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda using the provided $50$ expert demonstrations from RoboCasa.
+We evaluate our method and the baselines on one task of each atomic task category except navigation and group doors and drawers due to their similarity as shown in Fig.
+4: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda, and TurnOnStove using the provided $50$ expert demonstrations from RoboCasa.
+The prompts used were: "Mug", "Door", "Microwave button", "Faucet", "Soda can", "Knob" respectively.
 Baselines.
-We compare our method with three baseline methods in simulation: 1) BC-RNN: a behavior cloning method with recurrent neural network implementation; 2) BC-Transformer: a behavior cloning imitation learning method with transformer architecture as the policy implementation [54]; 3) Diffusion Policy: the image-based diffusion policy of [9].
-For real-world experiments, we compare our method against the diffusion policy.
+We compare our method with three baseline methods in simulation: 1) BC-RNN: a behavior cloning method with recurrent neural network implementation; 2) BC-Transformer: a common behavior cloning method with transformer architecture [43]; 3) Diffusion Policy: the image-based diffusion policy of [2]; for completeness we also compare against 4) S2-BC-Transformer: BC-Transformer trained with spatial-semantic data; and 5) GenDP: a multi-view RGB-D method [22].
+As RoboCasa does not provide depth, we compare against GenDP on InsertPencil and HangMug in SAPIEN  [44].
 In all our experiments, we train for $500$ epochs on an NVIDIA RTX 4090 GPU.
-We use a batch size of $64$ and a learning rate of $1\text{e}{-4}$.
+We set the initial learning rate to $1\text{e}{-4}$, applied a $500$-step linear warm-up, and then used a cosine learning rate scheduler for the remainder of training.
 Simulation Results.
-We take advantage of the open-sourced implementation for the baseline methods from RoboMimic [54].
-In the demonstration datasets for each task, we assume variations set by RoboCasa in the color and shape of the target object, as well as differences in the background environment.
-To test the performance of our method and the baselines, we evaluate $50$ trials for each task and the corresponding success rates are listed in Table I, demonstrating the superior performance of the proposed S2-Diffusion approach.
+For simulation experiments in RoboCasa, we take advantage of the open-sourced implementation for the baseline methods from RoboMimic [43].
+In the demonstration datasets for each task, we assume variations set by RoboCasa in the color of the target object, as well as differences in the background environment.
+We evaluate $100$ trials with the default seed for each task and the corresponding success rates and the $95\%$ confidence intervals (computed via the Wilson score method for binomial proportions) are listed in Table I.
+The results are demonstrating matching or statistically significantly superior performance of our S2-Diffusion approach.
 Due to the challenging variations in target object and background seen within the RoboCasa dataset, classical imitation learning policies struggle to solve the specific single-stage manipulation task.
-While the baseline methods—BC-RNN, BC-Transformer, and Diffusion Policy—show relatively poor performance, particularly on tasks like ServeMug, CloseDoor, and MoveSoda, S2-Diffusion consistently achieves the highest success rates.
-For instance, S2-Diffusion outperforms the other methods with a success rate of 0.72 on ServeMug and 0.78 on CloseDoor, while the best-performing baseline, BC-Transformer, achieves only 0.14 and 0.58, respectively.
-On other tasks like TurnOnMicrowave and MoveSoda, S2-Diffusion maintains a strong performance with success rates of 0.84 and 0.82.
-These results highlight the effectiveness of S2-Diffusion in improving task success across a diverse range of manipulation tasks.
-Similar to findings in RoboCasa [52], the image-based Diffusion Policy is sensitive to in object color and background.
+While the baselines—BC-RNN, BC-Transformer, and Diffusion Policy—show relatively poor performance, particularly on tasks like ServeMug and MoveSoda, S2-Diffusion consistently achieves the highest success rates.
+This lower performance of both BC and diffusion policy on RoboCasa is likely due to the high diversity of scenes in the expert training data.
+We also trained S2-BC-Transformer on the spatial-semantic dataset.
+We extended BC-Transformer because in our prior evaluation it exhibited comparable performance with DP.
+While S2-BC-Transformer achieves competitive results on some tasks, S2-diffusion demonstrates more robust performance.
+Additionally, S2-BC-Transformer outperforms the RGB-only version, highlighting the value of spatial-semantic representations for improving the generalization of imitation policies.
+Similar to findings in RoboCasa [41], the image-based DP is sensitive to in object color and background.
 In contrast, by leveraging spatial-semantic features our method shows better robustness and generality compared to state-of-the-art alternatives.
 We attribute this to the rich visual representation of combining semantic information and depth estimation for the workspace scene.
+Finally, Table II shows that our RGB method performs on par with the multi-view RGB-D method GenDP.
+Figure 5: Comparison of our S2-Diffusion and the baseline on two real-world environments: whiteboard wiping and bowl-to-bowl scooping.
+S2-Diffusion and the baseline are trained on red-whiteboard-wiping dataset and rice-bowl-to-bowl-scooping dataset respectively, then evaluated on the known instances and transferred to unseen instances of the two tasks.
+Note that for choco-cereal-btb-scooping, hearts-cereal-btb-scooping, mixed-cereal-btb-scooping, and green-whiteboard-wiping the baseline diffusion policy shows $0\%$ success rate.
 VI-B Real-World Experiments
@@
 A single camera providing RGB observation was mounted on the end-effector as well as task-specific tools - such as a sponge for the whiteboard wiping and a spoon for the scooping task.
-The teleoperation setup is shown in Fig.
-5.
 The language prompts for obtaining the semantic observations were "handwriting.
@@
 bowl." for the bowl-to-bowl scooping task.
-Figure 5: Real-world setup for the wiping task showing the teleoperation controller held by the operator on the right.
 Baselines and Ablations.
 We compare our S2-Diffusion method with the visuomotor diffusion method from [2] which uses RGB and proprioceptive information as a baseline, as well as perform a number of ablations on the scooping task - namely a version of our method that only has access to semantic observations (Semantic-Diffusion) and one that only observes the spatial observations (Spatial-Diffusion).
-Figure 6: Comparison of our S2-Diffusion and the baseline on two real-world environments: whiteboard wiping and bowl-to-bowl scooping.
-S2-Diffusion and the baseline are trained on red-whiteboard-wiping dataset and rice-bowl-to-bowl-scooping dataset respectively, then evaluated on the known instances and transferred to unseen instances of the two tasks.
-Note that for choco-cereal-btb-scooping, hearts-cereal-btb-scooping, mixed-cereal-btb-scooping, and green-whiteboard-wiping the baseline diffusion policy shows $0\%$ success rate.
+Finally, we evaluate generalization to transparent objects.
 Real-World Results.
 The real-world experimental results are reported in Fig.
-6.
+5.
 We evaluate our method and the baseline first on the seen environment (red-whiteboard-wiping and rice-bowl-to-bowl-scooping) and evaluate the transferability on other instances of the same category - namely black-whiteboard- and green-whiteboard-wiping and choco-cereal-, hearts-cereal-, mixed-cereal-bowl-to-bowl-scooping.
-Whiteboard-wiping: The task involves wiping scribbles of different colors (red, black, green) from a whiteboard using the end-effector-mounted wet sponge.
+Whiteboard-wiping.
+The task involves wiping scribbles of different colors (red, black, green) from a whiteboard using the end-effector-mounted wet sponge.
 The results are shown in Fig.
-6.
+5 including the $95\%$ Wilson confidence interval.
 We define success as the robot completely removing the targeted handwriting using a sponge within $15$ seconds.
@@
 These results highlight the ability of our S2-Diffusion to learn category-level skills for handling novel tasks without requiring additional training or fine-tuning, a capability that is essential for real-world robotic applications.
-Bowl-to-bowl-scooping: The scooping task is to scoop granular materials from one bowl to another.
+Note that the results of black-whiteboard-wiping are not statistically significant - we attribute this to the fact that the black observations are more similar to red than the green scribbles.
+Bowl-to-bowl-scooping.
+The scooping task is to scoop granular materials from one bowl to another.
 The amount of material successfully scooped into the target bowl is measured in grams.
@@
 We can see frames of the baseline diffusion policy and our method in Fig.
-7, where the baseline can only succeed on the instance-level task it was trained on and fails to scoop out cereal as the visual observations are too different.
+6, where the baseline can only succeed on the instance-level task it was trained on and fails to scoop out cereal as the visual observations are too different.
 Our method on the other hand extracts the semantic mask from the prompt as well as the spatial information via the synthetic depth observation.
@@
 The baseline can only succeed on the instance-level task it was trained on.
-Our method succeeds in learning a category-level skill trained on individual instances only using a single RGB observation.
+Our method succeeds in learning a category-level skill trained on individual instances.
 VI-C Ablation Evaluation
@@
 The results of these experiments are presented in Fig.
-8.
-The full model S2-Diffusion, which integrates both modalities, achieves the highest success rates, significantly outperforming the ablated versions both on the seen task and also the novel task.
+7.
+The full model S2-Diffusion, which integrates both modalities, achieves the highest success rates, significantly outperforming the Semantic-Diffusion and also etching out the Spatial-Diffusion versions both on the seen task and also the novel task.
 When the depth information is excluded, the performance of the policy drops drastically, with success rates falling to $0.4$ and $0.5$ for the seen and novel tasks, respectively.
 This significant decline underscores the critical role that spatial information plays in successfully executing 3D tasks.
-In contrast, when the mask (semantic representation) is removed, the performance also declines but to a lesser extent compared to the removal of depth information.
-While the policy’s ability to generalize across tasks is still reasonable in this case, it clearly benefits from the combined use of both modalities.
-From these results, we conclude that both semantic representation and depth estimation are indispensable for learning a generalizable skill trained on individual instances.
-The integration of spatial-semantic features enhances the robustness of the policy, enabling better performance across both familiar and novel environments, and ultimately leading to more effective task execution in 3D spaces.
+To evaluate generalization to transparent objects, we add a task where it has to close a transparent container.
+For reference, we also compare against a DP with depth camera input, trained directly on the transparent container.
+We find that S2 Diffusion only has a small reduction in success rate while the depth camera policy was not able to complete the task - illustrating the robustness of the S2 approach in common kitchen tasks involving transparent or reflective surfaces - where depth sensors struggle.
+In contrast, when the mask (semantic representation) is removed, the performance also declines, but to a lesser extent compared to the removal of depth information.
+While the policy’s ability to generalize across tasks is still reasonable in this case, it benefits from the combined use of both modalities.
 Figure 7: Ablation study of Semantic and Spatial-diffusion only on a seen task instance (left) and an unseen task instance (right).
-We show how the semantic component and spatial component contribute to the performance of the policy.
 S2-Diffusion outperforms models trained with only semantic or spatial observations.
@@
 As Fig.
-2 shows, the functional goal of flipping or scooping is discretized in different category-level skills such as bowl-to-bowl and pile-to-container scooping.
-While our method does generalize from instance-level tasks to category-level tasks, in its current form it is not able to generalize well from one instance-level skill to an instance of a different category task.
-i.e.
-when only training on rice-bowl-to-bowl the performance will degrade on a sand-pile-to-container task.
-While Fig.
-2 makes the ordering of skills into instance and category levels look straightforward, there are some corner cases that would need to be addressed if this view was expanded into a more complex ontology.
-For one, the functional goal can depend on more context than what is shown in our example.
-For instance, flipping is presented in our example as a way to turn over an object, but it could also be interpreted as an action the robot should perform with its own body.
-Furthermore, given how current learning-from-demonstration techniques are implemented, the category bowl-to-bowl scooping might require additional qualifiers to accurately encompass the correct instances.
-For example, if we were to generalize this skill to ice cream bowl-to-bowl scooping, the policy would fail, as the forces required are completely different from those observed in the training instances.
-Ice cream, especially when frozen, requires significantly more force to be scooped in the first place.
-However, as the ice cream begins to melt, and the required forces become more similar to those seen during training, there is a point at which our category skill bowl-to-bowl scooping would likely work.
-We leave exploring such nuances to future work.
-Another limitation of our work is its dependency on the performance of pre-trained vision foundation models.
-Naturally if the semantic or spatial estimate is poor, the resulting category-level skill may generalize poorly.
-To address this, we plan to investigate fine-tuning VLMs in conjunction with the downstream robot policy network in future research.
+2 illustrates, functional goals like flipping or scooping can be discretized into different category-level skills such as bowl-to-bowl and pile-to-container scooping.
+While our method generalizes from instance-level to category-level tasks, it does not generalize well across different categories; for example, a policy trained on rice-bowl-to-bowl scooping struggles with a sand-pile-to-container task.
+Although the abstraction hierarchy in Fig.
+2 appears straightforward, real-world functional goals can depend on more complex context and categories like bowl-to-bowl scooping may require additional qualifiers.
+For example, transferring frozen ice cream differs significantly from transferring cereals due to the forces required, although generalization may improve as the ice cream melts.
+Addressing such nuanced skill ontologies is an avenue for future work.
+Another limitation is the dependency on the performance of pretrained VLMs as well as proper text prompts for accurate semantic segmentation.
+Poor semantic or spatial estimates can degrade category-level skill generalization, particularly for greatly different object instances.
+Fine-tuning VLMs jointly with the policy network is a promising direction for mitigating this dependency.
 VIII CONCLUSION
-In this work, we have introduced S2-Diffusion, an open-vocabulary spatial-semantic diffusion policy that enables generalization from instance-level training data to category-level skills in robotic manipulation.
-Our method integrates semantic understanding with spatial representations, leveraging vision foundation models to learn a policy that is invariant to task-irrelevant visual changes and generalizes across different instances of the same category.
-This allows robots to transfer learned skills beyond their training data without requiring additional fine-tuning or retraining.
-Through extensive simulations and real-world evaluations, we demonstrated that S2-Diffusion outperforms the baselines. In particular we showed that:
-  • Spatial-Semantic representations enhance generalization – By combining spatial information with semantic segmentation, our method enables robots to focus on task-relevant features, improving instance-to-category transfer.
-  • Efficient real-time execution using only a single RGB camera – Unlike approaches that require multi-view setups or additional depth sensors, S2-Diffusion extracts meaningful scene representations from a single RGB image, making it practical for real-world deployment.
-  • Category-level generalization – Our evaluations on a number of robotic manipulation tasks, including real-world wiping and scooping, demonstrate that the method successfully generalizes across unseen instances within the same category, achieving high performance where baseline policies fail.
-In summary, S2-Diffusion policies are less constrained by their training distribution and constitute a step toward the ability of humans to generalize skills across variations in objects, materials, and environments instances.
+We introduced S2-Diffusion, an open-vocabulary spatial-semantic diffusion policy that enables generalization from instance-level training data to category-level skills in robotic manipulation.
+By integrating semantic understanding and spatial representations via vision foundation models, our method learns policies invariant to task-irrelevant visual changes, allowing skill transfer without additional fine-tuning.
+Through extensive simulations and real-world evaluations, we demonstrated that S2-Diffusion outperforms baselines.
+In particular:
+  • Spatial-Semantic representations enhance generalization: By combining spatial and semantic information, our method focuses on task-relevant features, enabling instance-to-category transfer.
+  • Efficient real-time execution: S2-Diffusion requires only a single RGB camera, avoiding multi-view setups or depth sensors, making it practical for real-world deployment and attractive for mobile settings.
+  • Category-level generalization: Our evaluations show strong generalization across unseen instances, achieving high performance where baseline policies fail.
+Overall, S2-Diffusion represents a step toward enabling robots to generalize skills across variations in objects, materials, and environments, similar to human capabilities.